OpenStudy (anonymous):

Which sets of numbers are closed under subtraction? Choose all answers that are correct. A. odd natural numbers B. rational numbers C. {0, 1} D. {0, 1, 2}

OpenStudy (anonymous):

@KonaFirestar @wio @dan815

OpenStudy (jtvatsim):

That's a fun one. Do you recall the definition of what it means to be closed under an operation?

OpenStudy (anonymous):

no

OpenStudy (jtvatsim):

So, to be "closed" basically means that when you perform the operation on the numbers in your set, your answer stays inside the set. Let me give an example.

OpenStudy (jtvatsim):

{0, 1, 2} is NOT closed under addition for example, because while 0 + 1 = 1 is in the set, 0 + 2 = 2 is in the set, and 1 + 1 = 2 is in the set, 2 + 1 = 3 is NOT in the set.

OpenStudy (jtvatsim):

Does that sort of make sense of what we mean by closed?

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

Alright, so let's try your question with subtraction

OpenStudy (jtvatsim):

Start with C and D, those will be easier to rule in or out :)

OpenStudy (jtvatsim):

So, is {0, 1} closed under subtraction?

OpenStudy (anonymous):

no

OpenStudy (jtvatsim):

Good, why not?

OpenStudy (anonymous):

because 1 is not in the set

OpenStudy (jtvatsim):

you mean -1 (negative 1) right? Because 0 - 1 = -1 not in the set.

OpenStudy (anonymous):

yes sry :p

OpenStudy (jtvatsim):

no probs :) just checking

OpenStudy (jtvatsim):

OK, cool, how about D?

OpenStudy (anonymous):

yes i think

OpenStudy (jtvatsim):

So, {0, 1, 2} is NOT closed... for the same reason 0 - 1 = -1 not in the set (still!) :)

OpenStudy (jtvatsim):

Make sense so far? Ready for A and B?

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

Alright, so are the odd natural numbers closed under subtraction?

OpenStudy (jtvatsim):

In other words, is the set {1, 3, 5, 7, 9, ...} closed under subtraction?

OpenStudy (jtvatsim):

"natural" means positive... just FYI :)